# vCache: Verified Semantic Prompt Caching

### v3

@@ Abstract
Semantic caches return cached LLM-generated responses for semantically similar prompts to reduce inference latency and cost. They embed cached prompts and store them alongside their response in a vector database. Embedding similarity metrics assign a numerical score to quantify the similarity between a request and its nearest neighbor prompt from the cache. Existing systems use the same static similarity threshold across all requests to determine whether two prompts can share similar responses. However, we observe that static thresholds do not give formal correctness guarantees, can result in unexpected error rates, and lead to suboptimal cache hit rates. This paper proposes vCache, the first verified semantic cache with user-defined error rate guarantees. It employs an online learning algorithm to estimate an optimal threshold for each cached prompt, enabling reliable cache responses without additional training. Our experiments show that vCache consistently meets the specified error bounds while outperforming state-of-the-art static-threshold and fine-tuned embedding baselines. We release the vCache implementation and benchmarks to support future research.
@@ 1 Introduction
Semantic caches can reduce large language model (LLM) inference latency by up to 100× by reusing previously generated responses for semantically similar prompts (Bang, 2023). Unlike exact-match caching, these systems can return cached responses not only for identical prompts but also for prompts that are similar in meaning, even if the input text differs (Razi et al., 2024). A semantic cache operates as follows. The cache $C$ embeds every prompt request $x$ into a vector $\mathcal{E}(x)\in\mathbb{R}^{d}$ and retrieves the semantically most similar prompt $nn(x)$, alongside its response $r(nn(x))$, from a vector database (Pan et al., 2024). The cache measures similarity (e.g., cosine similarity) between two embeddings using $s(x)=sim\bigl{(}\mathcal{E}(x),\mathcal{E}(y)\bigr{)}\in[0,1]$. If no sufficiently similar prompt is found, the LLM generates a response and adds its embedding along with the response to the vector database in the cache. In this paper, we study the reliability of semantic caches in returning correct responses for semantically similar prompts.
To determine whether a new prompt is sufficiently close to an existing prompt in the cache, state-of-the-art semantic caches rely on a user-selected threshold $t\in[0,1]$ (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025; Razi et al., 2024; Sudarsan and MasayaNishimaki, 2024). If $s(x)\geq t$, the system performs exploitation (cache hit) by returning the cached response $r(\mathrm{nn}(x))$. Otherwise, it performs exploration (cache miss) by querying the model for a new response $r(x)$. The cache adds $\mathrm{\mathcal{E}}(x)$ to the vector database, stores $r(x)$ in its metadata, and returns $r(x)$.
@@
We propose vCache, the first verified semantic cache with theoretical correctness guarantees. vCache learns a separate threshold (Figure 1) for each embedding in the cache, capturing the threshold variability observed in Figure 2. It requires no offline training, is agnostic to the underlying embedding model, and dynamically adapts its thresholds to the data distribution it encounters. As a consequence, vCache is robust to out-of-distribution inputs. To our knowledge, no prior work in semantic caching learns thresholds in an online manner.
We adopt a simple probabilistic framework to bound the error rate conditioned on a learned per-embedding threshold. When deploying vCache, the user specifies a maximum error rate bound, and the system maximizes the cache hit rate subject to this correctness constraint. Let $\mathbf{vCache}(x)$ denote the response returned by vCache. Let $\tau$ denote the exploration probability—a value monotonically decreasing in the likelihood of being correct—and calibrated so that the overall error rate remains below the user-specified bound (see Section 4). The decision rule modeling the probability of being correct for whether to exploit the cached response or explore an LLM inference is given by:
@@
We evaluate the effectiveness of vCache in terms of correctness guarantees and overall performance. To assess generalizability, we compare vCache across two embedding models, two LLMs, and five datasets. We find that vCache consistently meets the error-rate bounds and outperforms static threshold baselines, even when using fine-tuned embeddings. Specifically, it achieves up to 2×2$\times$2 × higher cache hit rates and reduces error rates by up to 6×6$\times$6 ×. Our main contributions are as follows:
@@
    We introduce three benchmarks that reflect classification-style, browser search queries, and open-ended prompt distributions.
@@
Figure 2: Results from 45k samples in the SemCacheClassification benchmark. Motivates the need for dynamic, embedding-specific thresholds. Top: Similarity values of correct and incorrect explorations exhibit highly overlapping distributions. Bottom: Optimal per-embedding thresholds vary substantially. See Appendix C.3 for details.
@@ 2 Related Work
LLM inference is both computationally expensive and latency-sensitive. To reduce these costs, semantic caches enable systems to return previously generated responses for not only identical, but also semantically similar prompts. Among existing systems, GPTCache is the most widely adopted open-source implementation, with over 7.5k GitHub stars (Bang, 2023). Several organizations and frameworks, including AWS (Razi et al., 2024), Microsoft (Dan Lepow, 2025), Redis-based caching in LiteLLM (lit, 2025), waLLMartCache (Dasgupta et al., 2025), and SCALM (Li et al., 2024), have adopted similar architectural patterns. A central limitation of existing systems is their reliance on a single global similarity threshold to determine whether a cached response should be returned to the user. This approach assumes that similarity correlates uniformly with correctness across all prompts and embeddings. However, as illustrated in Figure 2, the similarity distributions of correct and incorrect thresholds vary widely across prompts and embeddings. We describe the experimental setup underlying Figure 2 in Appendix C.3.
@@
Embedding Fine-tuning: One way to improve caching accuracy is to reshape the embedding space so that prompts yielding the same response are close together, while those with different responses are well-separated, increasing the reliability of threshold-based decisions. Zhu et al. (2024) propose a distillation-based method that fine-tunes embeddings to better align embeddings with equivalent responses. A related challenge arises in image retrieval, where systems must determine whether a nearest neighbor corresponds to the correct target class. Zhang et al. (2023a) address this by introducing the Threshold-Consistent Margin loss, which enforces tighter intra-class cohesion and clearer inter-class separation by selectively penalizing negative pairs. However, they require supervised training, are limited to open-source embedding models, and can fail to generalize to out-of-distribution data at inference time (Hajipour et al., 2024). vCache’s online learning algorithm does not require training, is model-agnostic, and generalized to out-of-distribution data (see SemCacheClassification benchmark in Figure 5).
Threshold Optimization: Threshold optimization attempts to learn a decision boundary over existing embeddings without modifying the embedding model itself (Zhang et al., 2023b). To our knowledge, no prior work in semantic caching learns thresholds online at inference time. Yet, as shown in Figure 2, the optimal similarity threshold varies significantly across embeddings, motivating embedding-specific and online threshold estimation. Related ideas have been explored in pen-set recognition and incremental learning. For example, Rudd et al. (2017) propose the Extreme Value Machine (EVM), which models class boundaries using extreme value theory to support generalization to unseen categories. However, these methods do not enforce user-defined error rate guarantees. In contrast, we introduce the first online algorithm that estimates per-embedding thresholds for semantic caches while satisfying a user-defined error bound.
@@
Even with high-quality embeddings and carefully tuned thresholds, semantic caches remain inherently approximate. Unless a threshold of 1.0 is used (effectively restricting cache hits to exact prompt matches), there is always a risk of returning incorrect responses (Razi et al., 2024; Jim Allen Wallace, 2024). Despite this, existing semantic caching systems rely on fixed thresholds to decide whether to return a cached response (Dasgupta et al., 2025; Li et al., 2024; Dan Lepow, 2025; Razi et al., 2024; lit, 2025; Bang, 2023). As a result, they offer no formal guarantees on accuracy or error rates, making it difficult to justify their reliability in production environments. To address this, we propose vCache, the first semantic caching system that combines competitive performance with a user-defined correctness guarantee, offering interpretable and controllable reliability.
@@ 3 Overview of Semantic Caching
$s(x)=sim(\mathcal{E}(x),\mathcal{E}(\mathrm{nn}(x)))\qquad c(x)=\begin{cases}1&\text{if }r\bigl{(}nn(x)\bigr{)}=r(x),\\ 0&\text{otherwise}.\end{cases}$ (4)
@@
vCache replaces static thresholding with an embedding-specific decision function that respects a user-defined maximum error rate bound $\delta$. If the function returns $\mathrm{exploit}$, the cache has enough certainty in the correctness of the response, exploits its knowledge, and returns the cached answer $r(\mathrm{nn}(x))$ (Algorithm 1, L5). If the function returns $\mathrm{explore}$, the cache has insufficient certainty and explores. Section 4 provides all the details of the vCache policy.
@@
Given a user-defined maximum error rate $\delta$, vCache maximizes the cache hit rate while ensuring the probability of correctness remains above $1-\delta$. Instead of relying on an unreliable static threshold or fine-tuned embeddings, vCache models probability distributions to make cache hit or miss decisions. The distributions are specific to each embedding in the cache and model the probability of correct cache hits for a given similarity value. To remain dataset agnostic and avoid costly offline training, vCache estimates these distributions online by selectively generating labels for uncertain similarity values. Since generating a label requires an LLM inference, equivalent in cost to not using a cache, vCache minimizes such inferences. We refer to such inferences as $\mathrm{explore}$ and classify them as a cache miss. For a given prompt $x$, the cached response $r(\mathrm{nn}(x))$ is considered uncertain when the observations $\mathcal{O}({nn(x))}$ do not provide sufficient evidence to determine whether the cached response is correct ($c(x)=1$). In contrast, if $\mathcal{O}({nn(x)})$ provides sufficient evidence, vCache proceeds with $\mathrm{exploit}$ by returning the cached response $r(\mathrm{nn}(x))$ without an LLM inference. The rest of this section formalizes these ideas and explains the vCache policy, $\mathbf{\mathcal{P}_{vCache}}$, in detail.
@@
1:$e_{x}\leftarrow\mathcal{E}(x)$
2:$y\leftarrow\mathrm{nn}(x)$
3:$s(x)\leftarrow sim(e_{x},\mathcal{E}(y))$
@@
vCache workflow for deciding whether to exploit a cached response (cache hit) or explore an LLM inference (cache miss). The decision relies on $\mathbf{\mathcal{P}_{vCache}}$ (Section 4) and respects the error rate guarantee $\delta$.
@@ 4.1 User Guarantee
Since vCache can only estimate $\mathbf{Pr}(c(x)=1\mid x,\mathcal{D})$ based on a limited number of samples, it accounts for the uncertainty in the estimation by considering the confidence band of $\mathbf{Pr}(c(x)=1\mid x,\mathcal{D})$. The modeling details and the vCache policy are presented in the next subsection.
@@
where $s(x)\in[0,1]$ is the similarity of $x$ with its near neighbour. $t\in[0,1]$ is an embedding-specific decision boundary parameter, and $\gamma>0$ is a parameter controlling the steepness of the function. The sigmoid form is well suited for this task: it induces a smooth and monotonic relationship between similarity and correctness probability and enables efficient maximum-likelihood estimation (MLE) of the threshold $t$ from labeled data. By fitting a continuous likelihood function rather than enforcing a hard threshold, vCache generalizes better from limited observations.
@@
Since these estimates are based on limited number of samples, estimating the true $\mathbf{Pr}(c(x){=}1|x,\mathcal{D})$, and thus the correct $\tau_{\mathrm{nn}(x)}(s(x))$ is not possible. To ensure we still achieve guarantees, vCache, instead computes a upper bound, say $\hat{\tau}$ for $\tau_{\mathrm{nn}(x)}(s(x))$ using pessimistic values for $t,\gamma$ from the $(1-\epsilon)$ confidence band for these points for some $\epsilon\in(0,1)$. Let these estimates be $t^{\prime}(\epsilon),\gamma^{\prime}(\epsilon)$. We compute $\hat{\tau}$ using,
@@
The details of why $\hat{\tau}\geq\tau_{\mathrm{nn}(x)}(s(x))$ and how to obtain confidence bands for $t$ and $\gamma$ are explained in Appendix B. Once we have $\hat{\tau}$, we have to ensure that the probability of exploration is above this value (see Eq 8 ). This is achieved by sampling a uniform random variable $u\sim\mathrm{Uniform}(0,1)$. If $u\leq\hat{\tau}$, the vCache explores, i.e., runs the LLM model to obtain the correct response. Otherwise, it exploits the cache by returning $r(\mathrm{nn}(x))$. This randomized policy ensures that, in expectation, the system explores sufficiently often to meet the correctness guarantee while maximizing cache usage when reliability is high. The exact algorithm of how $\mathrm{explore}$ and $\mathrm{exploit}$ decisions are made are presented in Algorithm 2.
@@
vCache makes the following assumptions,
  • The data $\mathcal{D}$ received by the cache is independently and identically drawn from the underlying distribution
  • The true probability of correctness of response match given similarity, i.e. $\mathbf{Pr}(c(x)=1|\mathcal{D},x)$ is well represented by the sigmoid family of functions (Eq 9).
Under these two assumptions, vCache policy ensures user-provided error-rate guarantees. This result is summarized in the theorem below.

### v4

@@ Abstract
Semantic caches return cached responses for semantically similar prompts to reduce LLM inference latency and cost. They embed cached prompts and store them alongside their response in a vector database. Embedding similarity metrics assign a numerical score to quantify the similarity between a request and its nearest neighbor prompt from the cache. Existing systems use the same static similarity threshold across all requests to determine whether two prompts can share similar responses. However, we observe that static thresholds do not give formal correctness guarantees, can result in unexpected error rates, and lead to suboptimal cache hit rates. This paper proposes vCache, the first verified semantic cache with user-defined error rate guarantees. It employs an online learning algorithm to estimate an optimal threshold for each cached prompt, enabling reliable cache responses without additional training. Our experiments show that vCache consistently meets the specified error bounds while outperforming state-of-the-art static-threshold and fine-tuned embedding baselines. We release the vCache implementation and three benchmarks to support future research.
@@ 1 Introduction
Large language models (LLMs) power applications ranging from conversational assistants to search engines and code generation, but their widespread use is limited by the high computational cost and inference latency (Zhao et al., 2023; Xiong et al., 2024; Achiam et al., 2023). Each new prompt requires multiple expensive forward passes through the model, which makes deployments costly and slow (Kwon et al., 2023). Prompt caching offers a natural way to mitigate this issue: if a prompt has already been answered, the system can return the cached response instead of performing another inference. Traditional exact string-match caching reduces cost by returning responses for repeated prompts, but it fails whenever the same intent is expressed in different words (Zhu et al., 2024). For example, a cache that already answered “Which city is Canada’s capital?” should also return the same response when later asked “What is the capital of Canada?”. Semantic caching addresses this limitation by retrieving responses for prompts that are semantically similar, even if their lexical form differs, and reduces inference latency by up to 100× (Bang, 2023). Semantic caches are effective in single-turn interactions with short to medium context, such as web search queries or classification tasks, where requests reappear in paraphrased forms but map to the same underlying response (Liu et al., 2024b; Wang et al., 2024). In this paper, we study the reliability of semantic caches in returning correct responses for semantically similar prompts.
Semantic caches operate as follows. The cache embeds every prompt request $x$ into a vector $\mathcal{E}(x)\in\mathbb{R}^{d}$ and retrieves the semantically most similar prompt $nn(x)$, alongside its response $r(nn(x))$, from a vector database (Pan et al., 2024). The cache measures similarity (e.g., cosine similarity) between two embeddings using $s(x)=sim\bigl(\mathcal{E}(x),\mathcal{E}(y)\bigr)\in[0,1]$. If no sufficiently similar prompt is found, an LLM generates a response and adds the embedded prompt along with the response to the vector database in the cache.
To determine whether a new prompt is sufficiently close to an existing prompt in the cache, state-of-the-art semantic caches rely on a user-selected threshold $t\in[0,1]$ (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025; Razi et al., 2024; Sudarsan & MasayaNishimaki, 2024). If $s(x)\geq t$, the system performs exploitation (cache hit) by returning the cached response $r(\mathrm{nn}(x))$. Otherwise, it performs exploration (cache miss) by querying the model for a new response $r(x)$. The cache adds $\mathrm{\mathcal{E}}(x)$ to the vector database, stores $r(x)$ in its metadata, and returns $r(x)$.
@@
We propose vCache, the first verified semantic cache with theoretical correctness guarantees. vCache learns a separate threshold (Figure 1) for each embedding in the cache, capturing the threshold variability observed in Figure 3. It requires no upfront training, is agnostic to the underlying embedding model, and dynamically adapts its thresholds to the data distribution it encounters. As a consequence, vCache is robust to out-of-distribution inputs. To our knowledge, no prior work in semantic caching 1) learns thresholds in an online manner and 2) guarantees their correctness.
We adopt a probabilistic framework to bound the error rate conditioned on a learned per-embedding threshold. When deploying vCache, the user specifies a maximum error rate bound $\delta$, and the system maximizes the cache hit rate subject to this correctness constraint (Figure 2). Let $\mathbf{vCache}(x)$ denote the response returned by vCache. Let $\tau$ denote the exploration probability—a value monotonically decreasing in the likelihood of being correct—and calibrated such that the overall error rate remains below the user‐specified bound (Section 4). The decision rule modeling the probability of being correct for whether to exploit the cached response or explore an LLM inference is given by:
@@
An illustrative overview of the vCache workflow and system architecture is provided in Appendix B.
We evaluate the effectiveness of vCache in terms of correctness guarantees and overall performance. To assess generalizability, we compare vCache across two embedding models, two LLMs, and five datasets. We find that vCache consistently meets the error-rate bounds and outperforms static threshold baselines, even when using fine-tuned embeddings. Specifically, it achieves up to $2\times$ higher cache hit rates and reduces error rates by up to $6\times$. Our main contributions are as follows:
@@
    We introduce three benchmarks, derived from five real-world datasets, that capture common semantic cache use cases: classification tasks, search queries, and open-ended prompt distributions.
@@ 2 Related Work
Existing semantic caches, such as GPTCache (Bang, 2023) and industry variants (Razi et al., 2024; Dan Lepow, 2025; lit, 2025; Dasgupta et al., 2025; Li et al., 2024), use a global similarity threshold to make cache hit decisions. This assumes a uniform correlation between similarity and correctness across all prompts and embeddings. However, as illustrated in Figure 3, similarity distributions vary widely, making fixed thresholds unreliable. Further details are provided in Appendix D.6.
@@
Embedding Fine-tuning: Zhu et al. (2024) propose a distillation-based method that fine-tunes embeddings for semantic caching, improving alignment between semantically equivalent prompts and their responses. A related challenge arises in image retrieval, where systems must determine whether a nearest neighbor corresponds to the correct target class. Zhang et al. (2023a) address this by introducing the Threshold-Consistent Margin loss, which enforces tighter intra-class cohesion and clearer inter-class separation by selectively penalizing negative pairs. However, they require supervised training, are limited to open-source embedding models, and can fail to generalize to out-of-distribution data at inference time (Hajipour et al., 2024). vCache’s online learning algorithm does not require training, is model-agnostic, and generalizes to out-of-distribution data (see Appendix E.1).
Threshold Optimization: Threshold optimization learns a decision boundary over existing embeddings without modifying the embedding model itself (Zhang et al., 2023b). To our knowledge, no prior work in semantic caching learns thresholds online at inference time. Yet, as shown in Figure 3, the optimal similarity threshold varies significantly across embeddings, motivating embedding-specific and online threshold estimation. Related ideas have been explored in incremental learning. For example, Rudd et al. (2017) propose the Extreme Value Machine (EVM), which models class boundaries using extreme value theory to support generalization to unseen categories. However, these methods do not guarantee user-defined error rates. In contrast, we introduce the first online algorithm that estimates per-embedding thresholds for semantic caches while satisfying a user-defined error bound.
@@
Even with high-quality embeddings and a presumably carefully tuned threshold, semantic caches remain inherently approximate. Unless a threshold of 1.0 is used (effectively restricting cache hits to exact prompt matches), there is always a risk of returning incorrect responses (Razi et al., 2024; Jim Allen Wallace, 2024). Despite this, existing semantic caching systems rely on fixed thresholds to decide whether to return a cached response (Dasgupta et al., 2025; Li et al., 2024; Dan Lepow, 2025; Razi et al., 2024; lit, 2025; Bang, 2023). As a result, they offer no formal guarantees on accuracy or error rates, making it difficult to justify their reliability in production environments. To address this, we propose vCache, the first semantic caching system that combines competitive performance with a user-defined correctness guarantee.
@@ 3 Overview of Semantic Caching
[…]
@@
vCache replaces static thresholding with an embedding-specific decision function that respects a user-defined error bound $\delta$. If the function returns $\mathrm{exploit}$, the cache is sufficiently confident and outputs the cached response $r(\mathrm{nn}(x))$ (Algorithm 1, L5). Otherwise, it returns $\mathrm{explore}$ by inferring the LLM. Section 4 provides further details.
Scope of Definitions
All quantities such as $nn(x)$, $s(x)$, and $c(x)$, are evaluated at a fixed (but arbitrary) point in time. Since the analysis is performed online, these definitions apply consistently across time steps. Policy discussion always refers to a specific embedding in the cache, with all parameters and estimates interpreted as conditional on it.
For ease of reference, a glossary of all symbols and functions is provided in Appendix A.
@@
Given a user-defined maximum error rate $\delta$, vCache maximizes the cache hit rate while ensuring the probability of correctness remains above $1-\delta$. Instead of relying on an unreliable static threshold or fine-tuned embeddings, vCache models probability distributions to make cache hit or miss decisions. The distributions are specific to each embedding in the cache and model the probability of correct cache hits for a given similarity value. To remain dataset agnostic and avoid costly offline training, vCache estimates these distributions online by selectively generating labels for uncertain similarity values. Since generating a label requires an LLM inference, equivalent in cost to not using a cache, vCache minimizes such inferences. We refer to such inferences as $\mathrm{explore}$ and classify them as a cache miss. For a given prompt $x$, the cached response $r(\mathrm{nn}(x))$ is considered uncertain when the observations $\mathcal{O}({nn(x))}$ do not provide sufficient evidence to determine whether the cached response is correct ($c(x)=1$). In contrast, if $\mathcal{O}({nn(x)})$ provides sufficient evidence, vCache proceeds with $\mathrm{exploit}$ by returning the cached response $r(\mathrm{nn}(x))$ without an LLM inference. The rest of this section formalizes these ideas and provides a detailed explanation of the vCache policy, $\mathbf{\mathcal{P}_{vCache}}$.
@@
[…]
@@
vCache workflow for deciding whether to exploit a cached response (cache hit) or explore an LLM inference (cache miss). The decision relies on the $\mathbf{\mathcal{P}_{vCache}}$ policy (Section 4) and guarantees a user-defined error rate bound $\delta$.
Figure 3: Results from 45k samples in the SemCacheClassification benchmark. Motivates the need for dynamic, embedding-specific thresholds. Top: Similarity values of correct and incorrect explorations exhibit highly overlapping distributions. Bottom: Optimal per-embedding thresholds vary substantially, indicating that no single threshold can suffice across embeddings (see Appendix D.6).
@@ 4.1 User Guarantee
Since vCache can only estimate $\mathbf{Pr}(c(x)=1\mid x,\mathcal{D})$ based on a limited number of samples, it accounts for the uncertainty in the estimation by considering the confidence band of $\mathbf{Pr}(c(x)=1\mid x,\mathcal{D})$. The modeling details and the vCache policy are presented in the following subsection.
@@
where $s(x)\in[0,1]$ is the similarity of $x$ with its near neighbour. $t\in[0,1]$ is an embedding-specific decision boundary parameter, and $\gamma>0$ is a parameter controlling the steepness of the function. The sigmoid form is well-suited for this task: it induces a smooth and monotonic relationship between similarity and correctness probability and enables efficient maximum-likelihood estimation (MLE) of the threshold $t$ from labeled data. By fitting a continuous likelihood function rather than enforcing a hard threshold, vCache generalizes better from limited observations.
@@
Since these estimates are based on a limited number of samples, estimating the true $\mathbf{Pr}(c(x){=}1|x,\mathcal{D})$, and thus the correct $\tau_{\mathrm{nn}(x)}(s(x))$ is not possible. To ensure we still achieve guarantees, vCache, instead computes a upper bound, say $\hat{\tau}$ for $\tau_{\mathrm{nn}(x)}(s(x))$ using pessimistic values for $t,\gamma$ from the $(1-\epsilon)$ confidence band for these points for some $\epsilon\in(0,1)$. Let these estimates be $t^{\prime}(\epsilon),\gamma^{\prime}(\epsilon)$. We compute $\hat{\tau}$ using,
@@
The details of why $\hat{\tau}\geq\tau_{\mathrm{nn}(x)}(s(x))$ and how to obtain confidence bands for $t$ and $\gamma$ are explained in Appendix C. Once we have $\hat{\tau}$, we have to ensure that the probability of exploration is above this value (see Eq 8 ). This is achieved by sampling a uniform random variable $u\sim\mathrm{Uniform}(0,1)$. If $u\leq\hat{\tau}$, the vCache explores, i.e., runs the LLM model to obtain the correct response. Otherwise, it exploits the cache by returning $r(\mathrm{nn}(x))$. This randomized policy ensures that, in expectation, the system explores sufficiently often to meet the correctness guarantee while maximizing cache usage when reliability is high. The exact algorithm of how $\mathrm{explore}$ and $\mathrm{exploit}$ decisions are made is presented in Algorithm 2. Figure 2 illustrates the vCache modeling, where each subplot shows one cached embedding. Green and red points indicate correct and incorrect responses to observed similarities. The yellow dashed curve is the sigmoid model, with threshold $\hat{t}$ obtained by MLE. The blue dashed curve represents the sigmoid fit based on confidence bounds, with threshold $t^{\prime}$ selected to meet the user-defined error bound $\delta$. The dark green curve $\tau(s)$ denotes the exploration probability, where for a given similarity $s$, vCache explores with probability $\tau(s)$ and exploits the cache otherwise.
@@
vCache makes two assumptions. First, the data $\mathcal{D}$ received by the cache is independently and identically drawn from the underlying distribution. Second, the true probability of correctness of response match given similarity, i.e. $\mathbf{Pr}(c(x)=1|\mathcal{D},x)$ is well represented by the sigmoid family of functions (Eq 9).. Under these assumptions, the vCache policy can provide user-defined error-rate guarantees, as summarized in the following theorem.